OpenStudy (anonymous):

Mary has twice as many pencils as she has pens. Tom borrows a pen. Now, Mary has ten less pens than pencils. How many pencils does Mary have?

OpenStudy (tkhunny):

C = Number of Pencils N = Number of Pens You're about half done. Let's see what you get.

OpenStudy (anonymous):

C= 2N

OpenStudy (anonymous):

So the total amount she has is 2n + N

OpenStudy (anonymous):

So now?

OpenStudy (tkhunny):

Okay, now translate the "Tom" information.

OpenStudy (anonymous):

2N + N = M so then when Tom takes the pen it becomes 2N + N = M is now 2N = M

OpenStudy (tkhunny):

We have: C = Number of Pencils N = Number of Pens Mary has twice as many pencils as she has pens. C = 2N Tom borrows a pen. Now, we've only N-1 Pens Now, Mary has ten less pens than pencils C - 10 = N-1 Translate EXACTLY what it says. No need to talk your way out of it.

OpenStudy (anonymous):

N= 9

OpenStudy (anonymous):

So C= 19 since there are ten less pens than pencils, right?

OpenStudy (anonymous):

Thank you for the help! I got it! :)

OpenStudy (tkhunny):

Woo-hoo!!!! Careful with that translating? Be as exact as you can.